The potential in a region between x = 0 and x = 6.00 m is V = a + bx, where a = 13.8 V and b = -4.30 V/m.

(a) Determine the potential at x = 0. ___V

Determine the potential at x = 3.00 m. ____V

Determine the potential at x = 6.00 m. _____V

(b) Determine the magnitude and direction of the electric field at x = 0.

magnitude ___V/m

direction ___

Determine the magnitude and direction of the electric field at x = 3.00 m.

magnitude___ V/m

direction____

Determine the magnitude and direction of the electric field at x = 6.00 m.

magnitude____ V/m

direction_____
Answer #1

given :

V = a + bx

a = 13.8

b = -4.30

part a )

at x = 0

V = a + b*0

V = a = 13.8 V

part b)

x = 3

V = 13.8 - 4.3 *3

V = 0.9 V

part c )

x = 6

V = 13.8 - 4.3 * 6

V = -12 V

part 2 )

E = -dV/dx

E = -b

E = 4.3 V/m

and E will be cosntant for every point

direction will be in +x direction